OpenStudy (anonymous):

Solve the triangle. A = 46°, a = 34, b = 27

OpenStudy (anonymous):

I got B = 34.8°, C = 99.2°, now I just need c

OpenStudy (jack1):



OpenStudy (jack1):

so Sine Rule: a/ sinA = b / sin B 34 / sin 46 = 27/ sin B sin B = 0.5712 B = 34.84 degrees Triangle Rule: A + B + C = 180 46 + 34.8 + C = 180 C = 99.2 so little c: can use sine rule again or cos rule for a change?

OpenStudy (anonymous):

It's okay..I found the answer.

OpenStudy (jack1):

sweet, sorry i took so long, was doing other stuff as well and getting distracted

OpenStudy (anonymous):

It's okay :)